OpenStudy (anonymous):

Find an equation using the cotangent function that has the same graph as y = 7 tan(x).

OpenStudy (jdoe0001):

\(\bf recall\implies cot(\theta)=\cfrac{1}{tan(\theta)}\implies tan(\theta)=\cfrac{1}{cot(\theta)}\) so... how could you change that function, with a trig identity that use cotangent instead?